If 1/4 of a gallon paint covers 1/8 of a fence, how much of the fence can be painted with 1 gallon of paint?
1
Expert's answer
2019-12-20T04:14:30-0500

To solve this, we have to first understand the fact that, if paint is more, it can cover more area of fence to color.

So, when (1/4) of a gallon covers (1/8) of fence, 1 gallon of color (¼ gallon of color < 1 gallon of color) will cover more area of fence for coloring.

So, when (1/4th) gallon covers 1/8th of a fence, 1 gallon covers

(1/8)/ (1/4) = (1/8) X (4) = 1/2 of a fence.

Answer: 1 gallon of color can cover 1/2 of the fence.